For the following unbalanced reaction, if 155g of Fe2O3 are reacted ,how many grams of Fe will you produce?


1
Expert's answer
2021-02-10T01:41:12-0500

The atomic masses of Fe and O are 55.8 g/mol and 16 g/mol respectively.
The number of moles is the ratio of mass to molar mass.
The number of moles of Fe =55.8g/mol
446g​=8 moles.
8 moles of Fe will give 4
8×2​=4 moles of Fe2​O3​
The molar mass of (Fe2​O3​) = 2(55.8)+3(16)=159.6 g/mol
The mass of Fe2​O3​ obtained =159.6×4=640 g